# The HITRAN2024 methane update

### v1

@@ 4 Conclusion
The parameters of more than 58,000 lines of 12CH4, 13CH4, and 12CH3D have been updated for HITRAN20204 using the new line lists provided. Transitions wavenumbers, intensities, and lower-state energy levels were taken directly from these lists to replace the values from HITRAN2020. Laboratory spectra in the pentad and octad regions were used by the HITRAN team to validate and choose between the overlapping data sets, while other regions used the validations from the other groups. 40,000 lines of the various isotopologues have been added, including some lower-intensity lines that were previously missing. In regions that were previously empirically determined, this work allowed us to reassign some of the intensities to the correct isotopologues. Assignments of many lines have been completed between the octad and triacontad regions.
@@ Appendix A Abbreviations
    TheoReTS – Theoretical Reims-Tomsk Spectral data

### v2

@@ 4 Conclusion
The parameters of more than 80,000 lines of 12CH4, 13CH4, and 12CH3D have been updated for HITRAN20204 using the new line lists provided. Transitions wavenumbers, intensities, and lower-state energy levels were taken directly from these lists to replace the values from HITRAN2020. Laboratory spectra in the pentad, octad and tetradecad regions were used by the HITRAN team to validate and choose between the overlapping data sets, while other regions relied on the validations from the other groups. We observed a great drop in the residuals of the octad and tetradecad regions spectra particularly thanks to the atm.161 list and the new data from Reed et al.[reed_multilaboratory_2025] and Yan et al.. The effects on the pentad region were limited, but still showed improvements, especially at high er energies. 45,000 lines of the various isotopologues have been added, including some lower-intensity CH3D lines that should have been in the previous update but were missing, and increasing the region covered from 12,000 to 14,000 cm-1 with experimental data. In regions that were previously empirically determined, this work allowed us to reassign some of the intensities to the correct isotopologues. Assignments of many lines have been completed between the octad and triacontad regions.
@@ Appendix A Abbreviations
    TheoReTS – Theoretical Reims-Tomsk Spectral data computed from potential energy and dipole moment surfaces